OpenStudy (anonymous):

a box is pushed 40 m by a mover. the amount of work done was 2,240j. how much force was exerted on the box a box is pushed 40 m by a mover. the amount of work done was 2,240j. how much force was exerted on the box @Physics

OpenStudy (stormfire1):

\[W = F * distance * cos \theta\]Where theta is the angle between the force and the displacement vectors.\[2240 = F * 40m * cos(0)\]\[cos(0) = 1\]\[2240 = F * 40m\]\[F = 56N\]

OpenStudy (anonymous):

if the W=2.240J then then F=0.056N using the same processure from above Stormfires work